OpenStudy (anonymous):

In a geometric sequence, a4=7 and a6=28. Use the geometric mean to find the value of a5

OpenStudy (cwrw238):

if r is common ratio r^2 = a6/a4 = 28/7 = 4 need any more help?

OpenStudy (cwrw238):

actually i'm not using the geometric mean to do this!! geometric mean of a and b is sqrt(ab) so if you plug in a = 7 and b = 28 you will get the value of a5

OpenStudy (cwrw238):

a5 = sqrt ( 7 *28)